# DeblurDiNAT: A Lightweight and Effective Transformer for Image Deblurring

### v1

@@ 3.3 Divide and Multiply Feed-Forward Network
Figure 2: The structures of (a) Transformer block, (b) local channel modulation self-attention (CMSA) with a dilation factor of $1$, (c) global CMSA with a dilation factor of $\frac{n}{k}$, (d) cross-channel learner (CCL), and (e) divide and multiply feed-forward network (DMFN). The kernel sizes used in 1D Conv and Depthwise Conv are set to $3$.
Our proposed DMFN, as shown in Fig. 2 (e), incorporates a $1\times 1$ point-wise convolution layer followed by a $3\times 3$ depth-wise separate convolution layer to explicitly model channel interactions in the features. Using depth-wise convolution is inspired by Restormer and FFTformer [47, 23], where GELU functions are also introduced as part of gating mechanisms for controlled information flow. We argue that this non-linear function not only increases computational costs, but also hinders the deblurring performance of Transformers. Our divide and multiply strategy is simple while effective. From a layer normalized feature $X\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, DMFN first applies point-wise convolutions to expand the feature channels by a factor $\gamma=2$. After that, $3\times 3$ depth-wise convolutions are utilized to aggregate spatial context, generating $X_{0}\in\mathbb{R}^{\hat{H}\times\hat{W}\times 2\hat{C}}$. Next, we equally divide the intermediate feature map along the channel dimension, yielding $X_{1}$ and $X_{2}$, so that their element-wise multiplication result is of shape $\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$. The overall DMFN process is formulated as:
$\hat{X}=Conv_{1\times 1}(DM(X))+X,$ (6)
$DM(X)=X_{1}\odot X_{2},$
where $Conv_{1\times 1}$ denotes a convolution with filter size of $1\times 1$; $DM$ represents our divide and multiply strategy; $X_{1}$ and $X_{2}$ are the two equally divided feature maps of size $\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$.
3.4 Lightweight Gated Feature Fusion
Figure 3: The structures of (a) multi-scale AFF [5], (b) multi-scale LGFF, (c) gated feed-forward network (GFN), (d) same-scale simplified feature fusion (SFF), (e) same-scale simplified gated feature fusion (SGFF), (f) same-scale LGFF.
Most feature fusion methods in deblurring Transformers only aggregate same-scale data, which fail to exploit the full potential of feature hierarchies in images. We propose lightweight gated feature fusion (LGFF) that is also able to merge multi-scale features, illustrated in Fig. 3 (b)-(c). LGFF is formulated as:
$LGFF_{1}^{out}=LGFF_{1}(EB_{1}^{out},(EB_{2}^{out})^{\times 2},(EB_{3}^{out})^{\times 4}),$ (7)
$LGFF_{2}^{out}=LGFF_{2}((EB_{1}^{out})^{\times\frac{1}{2}},EB_{2}^{out},(EB_{3}^{out})^{\times 2}),$
@@ 3.4 Lightweight Gated Feature Fusion
$feature^{out}=LGFF(feature_{1}^{in},feature_{2}^{in}),$ (8)
where $feature_{1}^{in}$ and $feature_{2}^{in}$ are of the same shape. See Fig. 3 (f).
Different from the multi-scale asymmetric feature fusion (AFF) adopted in MIMO-Unet [5], we take advantage of depth-wise convolution and point-wise convolution to preprocess the concatenated features and reduce the number of channels. After that, a feed-forward network with two parallel paths is utilized to control the information flow across multiple depth levels.
@@ 4.1 Datasets and Implementation Details
We propose DeblurDiNAT-S (Small) and DeblurDiNAT-L (Large), the two variants of our model. The number of Transformer blocks in DeblurDiNAT-S are $[4,6,8]$ at each level, attention heads are $[2,4,8]$, and the number of channels are $[64,128,256]$. The refinement stage contains 4 Transformer blocks. For DeblurDiNAT-L, one more residual block is appended to each encoder; the number of Transformer blocks increment to $[6,12,18]$. We trained both models using Adam optimizer with the initial learning rate of $10^{-4}$ which was gradually decayed to $10^{-7}$ by the cosine annealing strategy. The batch size was set to $8$. We augmented input data using random cropping, flipping and rotation. We pretrained DeblurDiNAT models for 3000 epochs with a patch size of $256\times 256$ and further trained them for another 1000 epochs with a patch size of $512\times 512$. DeblurDiNAT was trained on eight NVIDIA A100 GPUs.
@@ 4.2.1 Quantitative Analysis
We compare DeblurDiNAT with existing SOTA deblurring models, including CNNs [33, 25, 38, 11, 48, 5, 32] and Transformers [47, 42, 40, 23]. All models were only trained on the GoPro dataset and tested on the GoPro [33], HIDE [36], RealBlur-R [35] and RealBlur-J [35] datasets. The image quality metrics PSNR/SSIM, and inference time are tested with full resolution $1280\times 720$ images on a single NVDIA A100 GPU. The number of FLOPs and GPU consumption are calculated with an input tensor of size $256\times 256$ on an NVIDIA 3090 GPU. The GPU consumption is computed by the “torch.cuda.max memory allocated” function.
@@
Tab. 1 shows that our proposed DeblurDiNAT-L holds the second position in the PSNR score rankings, following FFTformer [23], on the two synthetic datasets GoPro [33] and HIDE [36]. Tab. 2 compares the efficiency of our model to existing deblurring Transformers [47, 42, 40, 23]. It reports that DeblurDiNAT-L has $61\%$ less GPU memory consumption and runs 3.21×3.21$\times$3.21 × faster than FFTformer [23] which heavily involves complex matrix operations. Qualitative analysis (see Sec. 4.2.2) presents that our restored images are visually closer to the clean images than those of FFTformer [23]. DeblurDiNAT-S achieves satisfactory deblurring quality, which is comparable to the best CNN-based deblurring model MPRNet [32], while with $87\%$ fewer FLOPs and $53\%$ less inference time, as shown in Tab. 2. The PSNR scores of DeblurDiNAT-L measured on the RealBlur-J and RealBlur-R datasets [35] are higher than those of CNNs. Compared to Uformer-B [42] with the highest PSNR on RealBlur [35], our DeblurDiNAT-L also demonstrates competitive image quality scores and runs 4.49×4.49$\times$4.49 × faster with $68\%$ fewer parameters. Importantly, Tab. 1 reports that DeblurDiNAT-L obtains an average boost of $0.08$ dB over the previous best approach FFTformer [23], which proves that our method demonstrates the strongest generalization ability. From Tab. 1 and Tab. 2, we argue that our DeblurDiNAT guarantees efficacy, efficiency and robustness quantitatively.
@@
Table 1: Quantitative evaluations on the GoPro [33], HIDE [36], RealBlur-R [35] and RealBlur-J [35] datasets. The best, second-best and third-best PSNR scores are highlighted, underlined and double-underlined. The average PSNR is calculated over the four datasets with equal weights to measure model robustness quantitatively.
GoPro HIDE RealBlur-R RealBlur-J Avg.
Method PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR
Nah et al. [33] 29.08 0.914 25.73 0.874 32.51 0.841 27.87 0.827 28.80
DeblurGAN-v2 [25] 29.55 0.934 26.61 0.875 35.26 0.944 28.70 0.866 30.03
SRN [38] 30.26 0.934 28.36 0.915 35.66 0.947 28.56 0.867 30.71
DSD [11] 30.90 0.935 29.11 0.913 - - - - -
DMPHN [48] 31.20 0.940 29.09 0.924 35.70 0.948 28.42 0.860 31.10
MIMO-Unet+ [5] 32.45 0.957 29.99 0.930 35.54 0.947 27.63 0.837 31.40
MPRNet [32] 32.66 0.959 30.96 0.939 35.99 0.952 28.70 0.873 32.08
Restormer [47] 32.92 0.961 31.22 0.942 36.19 0.957 28.96 0.879 32.32
Uformer-B [42] 33.06 0.967 30.90 0.953 36.19 0.956 29.09 0.886 32.31
Stripformer [40] 33.08 0.962 31.03 0.940 36.08 0.954 28.82 0.876 32.25
FFTformer [23] 34.21 0.969 31.62 0.946 35.87 0.953 27.75 0.853 32.36
DeblurDiNAT-S 32.85 0.961 30.65 0.936 35.92 0.954 28.80 0.877 32.06
DeblurDiNAT-L 33.42 0.965 31.28 0.943 36.07 0.956 28.99 0.885 32.44
@@
DeblurDiNAT-S 9.1 99.46 21.7 4.7
DeblurDiNAT-L 16.1 152.61 150.5 7.4
4.2.2 Qualitative Analysis
The PSNR/SSIM metrics may not precisely measure the visual quality of deblurred images as perceived by human eyes. In this section, we present that the deblurred images produced by DeblurDiNAT-L stand out for their superior visual quality on both synthetic and real-world datasets [33, 36, 35]. All deblurring models are trained only on GoPro [33]. PSNR values are calculated based on the full-size images. Fig. 4 and Fig. 5 show that our method produces cleaner plate numbers and vivid facial details. Comparing a vast number of deblurred images using human eyes, we also notice that our DeblurDiNAT-L demonstrates a stronger generalization ability over existing models. For example, Fig. 6 shows that FFTformer [23] with the highest PSNR score tested on GoPro introduces unpleasant color changes when applied to the RealBlur-J dataset [35]; Uformer-B [42] with the best quantitative results on RealBlur datasets [35] might distort the appearance of texts according to our experiments. However, ours not only successfully preserves original picture colors but also clearly restores local details and structures of objects with real-world blur artifacts. Fig. 7 shows that DeblurDiNAT-L generates tidy texts and colorful graphic symbols from a blurry image in low-light environments, where the other networks fall short (feel free to zoom images in and out). From the above comparisons, we assert that DeblurDiNAT-L generates deblurred images with the best overall visual quality. More visual results are provided in supplementary materials.
@@ 4.3 Ablation Study
In this section, we conduct ablation experiments to evaluate each component of DeblurDiNAT on the GoPro dataset [33]. We trained all models on two NVDIA A100 GPUs with a patch size of $128\times 128$ for 3000 epochs only. The test settings are the same as the ones in Sec. 4.2.1.
4.3.1 Effect of ADFS
@@
Local or global PSNR/SSIM Params (M) FLOPs (G) Time (ms)
only local 31.73/0.951 9.082 99.41 19.6
only global 31.96/0.953 9.082 99.41 20.3
local$+$global 31.98///0.953 9.082 99.41 19.7
Alternating Dilation Factor Structure (ADFS) is essentially a global-local attention strategy. CMSA enhances the learning of more local representations by decreasing the dilation factor, allowing to capture region-specific blurred patterns. Using local CMSA and global CMSA alternately obtains gains of $0.25$ dB over the network with only local CMSA and $0.02$ dB compared to the architecture with only global CMSA (see Tab. 3). It is worth mentioning that such an alternating method adds very few computational costs and only increases the inference time by $0.51\%$.
@@
Network Component PSNR/SSIM Params (M) FLOPs (G)
Feature fusion AFF [5]$+$SFF $31.88/0.953$ 9.123 100.31
LGFF(SA$=$True)$+$SFF $31.88/0.953$ 9.096 99.71
LGFF(SA$=$True)$+$GFN $31.89/0.953$ 9.079 99.28
LGFF$+$GFN $31.89/0.953$ 8.995 97.13
LGFF$+$LGFF 31.96///0.953 9.082 99.41
Table 5: Influence of CCL and DMFN. In the CCL group, the positions of CCL and impacts of different global pooling methods are investigated.
@@
CCL CMSA w/o CCL $31.98/0.953$ 99.41 19.7
CCL before SA$+$GAP $31.99/0.954$ 99.46 22.6
CCL after SA$+$GAP $32.01/0.954$ 99.46 22.2
@@
4.3.2 Effect of LGFF
We design LGFF as a multi-scale feature fusion module, while further experiments in Tab. 4 show that it is also suitable for same-scale features. Fig. 3 compares the structures of several fusion units. In Tab. 4, replacing AFF [5] by LGFF (enhanced with SA) yields the same quality scores with $27$ thousands fewer parameters and $0.6$ G fewer FLOPs. Removing the SA layers from LGFF does no harm to PSNR and SSIM results, and brings about greater model efficiency, as shown in the third and fourth rows of Tab. 4. It also presents that applying LGFF to both multi-scale and same-scale feature fusion provides an overall gain of $0.08$ dB and improved efficiency over the baseline.
4.3.3 Effect of CCL
We investigate how different positions of cross-channel learner (CCL) in CMSA and various pooling strategies used in CCL affect deblurring performance. Tab. 5 shows that, our parallel design obtains a $0.08$ dB gain and $2.26\%$ speedup over the strategy placing CCL before the self-attention blocks. The global average pooling (GAP) in CCL leads to $0.18$ dB boost over the global max pooling (GMP). Compared to the self-attention without channel modulation, our proposed CMSA improves PSNR by $0.09$ dB.
4.3.4 Effect of DMFN
Experiment data in Tab. 5 demonstrates that our straightforward DMFN brings a performance improvement of $0.07$ dB PSNR compared to GDFN [47] with gating mechanisms. Although both of them share similar numbers of FLOPs, GDFN offers a $1.5\%$ speedup.
4.3.5 Analysis on Self Attention
Table 6: Comparison experiments among different self-attention mechanisms.
@@ 4.3.5 Analysis on Self Attention
CMSA 32.07 9.082 99.46 22.1 4.73
We compare CMSA with self-attention mechanisms used in Restormer [47], Stripformer [40] and FFTformer [23], three SOTA Transformers designed for image deblurring. We replace the self-attention modules CMSA in DeblurDiNAT with theirs. As reported in Tab. 6, ours outperforms the other attention methods in terms of PSNR and GPU memory consumption. Compared to SA employed in FFTformer [23], ours achieves $0.03$ dB higher PSNR, 2.45×2.45$\times$2.45 × faster speed and $49\%$ less GPU memory consumption.

### v2

@@ 3.3 Divide and Multiply Feed-Forward Network
Figure 2: The structures of (a) Transformer block, (b) local CMSA with a dilation factor of $1$, (c) global CMSA with a dilation factor of $\frac{n}{k}$, (d) LCCL, and (e) DMFN. The kernel sizes used in 1D Conv and Depthwise Conv are set to $3$.
As shown in Fig. 2 (e), DMFN removes the activation function from regular feed-forward networks, motivated by the simple gate [3]. We claim that DeblurDiNAT is still a powerful non-linear model, because Leaky ReLU in the encoder path ensures that the extracted feature maps contain non-linearity, and feature fusion modules with GELU propagate features with non-linearity across the network. From a layer normalized feature $\mathbf{X}\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$, DMFN first applies point-wise convolutions to expand the feature channels by a factor $\gamma=2$. After that, $3\times 3$ depth-wise convolutions are utilized to aggregate spatial context, generating $\mathbf{X}_{0}\in\mathbb{R}^{\hat{H}\times\hat{W}\times 2\hat{C}}$. Next, the intermediate feature map $\mathbf{X}_{0}$ is equally divided along the channel dimension, yielding $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, so that their element-wise multiplication result is of shape $\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}$. The overall process of DMFN with normalization is formulated as:
$\hat{\mathbf{X}}$ $=W_{p}^{0}\mathrm{DM}(\mathbf{X})+\mathbf{X},$ (5)
$\mathrm{DM}(\mathbf{X})$ $=W_{d}^{1}W_{p}^{1}\mathcal{L}(\mathbf{X})\odot W_{d}^{2}W_{p}^{2}\mathcal{L}(\mathbf{X}),$
where $W_{p}^{(\cdot)}$ and $W_{d}^{(\cdot)}$ denote the convolution with filter size of $1\times 1$ and $3\times 3$; $\mathrm{DM}$ represents our divide and multiply strategy.
3.4 Lightweight Dual-Stage Feature Fusion
Figure 3: Structures of (a) multi-scale LDFF, (b) same-scale LDFF, (c) ECR, (d) CFM, (e) multi-scale AFF [6], (f) same-scale simplified gated feature fusion (SGFF). Same-scale simplified feature fusion (SFF) is a regular convolutional layer.
We propose two variants of lightweight dual-stage feature fusion (LDFF), the multi-scale one and the same-scale one, as shown in Fig. 3 (a)-(b). As a preprocessing step, LDFF first resizes input features if they are from different encoder levels, and concatenate them along the channel dimension. The main body of LDFF is composed of two stages, efficient channel reduction (ECR) and complementary feature mixing (CFM). In ECR, a $1\times 1$ convolution is used to merge cross-channel contextual information, and a $3\times 3$ depth-wise convolution is for local content aggregation in the spatial domain. Then, CFM splits a normalized feature map into two complementary features by linear layers, passing through two parallel branches, resulting in a linear feature in the identity path and a non-linear one in the GELU gated path. The two complementary features are merged by element-wise multiplications and additional linear projection layers to refine image details. The two multi-scale LDFF blocks in DeblurDiNAT are formulated as:
$\mathrm{LDFF}_{1}^{out}=\mathrm{LDFF}_{1}(\mathrm{EB}_{1}^{out},(\mathrm{EB}_{2}^{out})^{\times 2},(\mathrm{EB}_{3}^{out})^{\times 4}),$ (6)
$\mathrm{LDFF}_{2}^{out}=\mathrm{LDFF}_{2}((\mathrm{EB}_{1}^{out})^{\times\frac{1}{2}},\mathrm{EB}_{2}^{out},(\mathrm{EB}_{3}^{out})^{\times 2}),$
@@ 4.1 Datasets and Implementation Details
We propose DeblurDiNAT-S (Small) and DeblurDiNAT-L (Large), the two variants of our model. The number of Transformer blocks in DeblurDiNAT-S are $[4,6,8]$ at each level, attention heads are $[2,4,8]$, and the number of channels are $[64,128,256]$. The refinement stage contains 4 Transformer blocks. For DeblurDiNAT-L, one more residual block is appended to each encoder; the number of Transformer blocks increment to $[6,12,18]$. We trained both models using Adam optimizer with the initial learning rate of $2\times 10^{-4}$ which was gradually decayed to $10^{-7}$ by the cosine annealing strategy. The batch size was set to $8$. We augmented input data using random cropping, flipping and rotation. We pretrained DeblurDiNAT models for 4000 epochs with a patch size of $256\times 256$ and further trained them for another 2000 epochs with a patch size of $512\times 512$. DeblurDiNAT was trained on eight NVIDIA A100 GPUs.
@@ 4.2 Performance Comparison
Quantitative Analysis. We compare DeblurDiNAT with existing SOTA deblurring models, including CNNs [32, 26, 37, 12, 47, 6, 46] and Transformers [45, 41, 38, 24]. All models were only trained on the GoPro dataset and tested on the GoPro [32], HIDE [35], RealBlur-R [34] and RealBlur-J [34] datasets. The image quality metrics PSNR/SSIM, and inference time are tested with full resolution $1280\times 720$ images on a single NVDIA A100 GPU. The number of FLOPs and GPU consumption are calculated with an input tensor of size $256\times 256$ on an NVIDIA 3090 GPU. The GPU consumption is computed by the “torch.cuda.max memory allocated” function.
@@
Tab. 1 reports that DeblurDiNAT-L obtains an average boost of $0.18$ dB over the previous SOTA approach FFTformer [24]. More specifically, Tab. 1 shows that DeblurDiNAT-L holds the second position in all the PSNR score rankings on synthetic datasets [32, 35] and real-world datasets [34]. It proves that our robust DeblurDiNAT-L demonstrates the strongest generalization ability. Tab. 2 compares the efficiency of our model to existing deblurring Transformers [45, 41, 38, 24]. It reports that DeblurDiNAT-L has $57\%$ less GPU memory consumption and runs 2.65×2.65$\times$2.65 × faster than FFTformer [24] which heavily involves complex matrix operations. The PSNR scores of DeblurDiNAT-L measured on the unknown RealBlur datasets [34] are higher than those of all models except Uformer-B [41]. Compared to Uformer-B [41], our DeblurDiNAT-L obtains an average gain of $0.23$ dB and runs 3.70×3.70$\times$3.70 × faster with $68\%$ fewer parameters. DeblurDiNAT-S outperforms nearly all models but FFTformer [24] on the synthetic GoPro dataset [32]. Compared to nearest competitors [6, 46, 45, 41, 38], DeblurDiNAT-S has $43.48\%$-$82.12\%$ fewer parameters. From quantitative comparisons Tab. 1 and Tab. 2, we argue that our DeblurDiNAT guarantees effectiveness, efficiency and robustness. Qualitative analysis (see Sec. 4.2) presents that our restored images are visually closest to the clean images.
@@
Table 1: The average PSNR is calculated over the GoPro [32], HIDE [35], RealBlur-R [34] and RealBlur-J [34] datasets with equal weights to measure the generalization ability of each deblurring model. The best, second-best and third-best PSNR scores are highlighted, underlined and double-underlined.
Avg. GoPro HIDE RealBlur-R RealBlur-J
Method PSNR PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Nah et al. [32] 28.80 29.08 0.914 25.73 0.874 32.51 0.841 27.87 0.827
DeblurGAN-v2 [26] 30.03 29.55 0.934 26.61 0.875 35.26 0.944 28.70 0.866
SRN [37] 30.71 30.26 0.934 28.36 0.915 35.66 0.947 28.56 0.867
DSD [12] - 30.90 0.935 29.11 0.913 - - - -
DMPHN [47] 31.10 31.20 0.940 29.09 0.924 35.70 0.948 28.42 0.860
MIMO-Unet+ [6] 31.40 32.45 0.957 29.99 0.930 35.54 0.947 27.63 0.837
MPRNet [46] 32.08 32.66 0.959 30.96 0.939 35.99 0.952 28.70 0.873
Restormer [45] 32.32 32.92 0.961 31.22 0.942 36.19 0.957 28.96 0.879
Uformer-B [41] 32.31 33.06 0.967 30.90 0.953 36.19 0.956 29.09 0.886
Stripformer [38] 32.25 33.08 0.962 31.03 0.940 36.08 0.954 28.82 0.876
FFTformer [24] 32.36 34.21 0.969 31.62 0.946 35.87 0.953 27.75 0.853
DeblurDiNAT-S 32.18 33.15 0.963 30.92 0.939 35.93 0.953 28.73 0.874
DeblurDiNAT-L 32.54 33.63 0.967 31.47 0.944 36.09 0.955 28.98 0.885
@@
DeblurDiNAT-S 9.1 99.46 24.3 5.2
DeblurDiNAT-L 16.1 152.61 182.8 8.2
Qualitative Analysis. The PSNR/SSIM metrics may not precisely measure the visual quality of deblurred images as perceived by human eyes. In this section, we present that the deblurred images produced by DeblurDiNAT-L stand out for their superior visual quality on both synthetic and real-world datasets [32, 35, 34]. All deblurring models are trained only on GoPro [32]. PSNR values are calculated based on the full-size images. Fig. 4 and Fig. 5 show that our method produces cleaner plate numbers and vivid facial details. Comparing a vast number of deblurred images using human eyes, we also notice that our DeblurDiNAT-L demonstrates a stronger generalization ability over existing models. For example, Fig. 6 shows that FFTformer [24] with the highest PSNR score tested on GoPro introduces unpleasant color changes when applied to the RealBlur-J dataset [34]; Uformer-B [41] with the best quantitative results on RealBlur datasets [34] might distort the appearance of texts according to our experiments. However, ours not only successfully preserves original picture colors but also clearly restores local details and structures of objects with real-world blur artifacts. Fig. 7 shows that DeblurDiNAT-L generates tidy texts and colorful graphic symbols from a blurred image in low-light environments, where the other networks fall short (feel free to zoom images in and out). From the above comparisons, we assert that DeblurDiNAT-L generates deblurred images with the best overall visual quality. More visual results are provided in supplementary materials.
@@ 4.3 Ablation Study
In this section, we conduct ablation experiments to evaluate each component of DeblurDiNAT on the GoPro dataset [32]. We trained all models on two NVDIA A100 GPUs with a patch size of $128\times 128$ for 3000 epochs only.
Effect of ADFS.
@@
Local or global PSNR/SSIM Time (ms) Params (M) FLOPs (G)
only local 31.73/0.951 19.6 9.082 99.41
only global 31.96/0.953 20.3 9.082 99.41
local+global 31.98///0.953 19.7 9.082 99.41
(a) Reference
(b) Blurred
(c) Only global
(d) Local+global
Figure 8: Effects of local attention on learning localized features.
Alternating Dilation Factor Structure (ADFS) is essentially a hybrid global-local attention strategy. As shown in Tab. 3, although DeblurDiNAT with hybrid attention and only global attention achieve similar metric scores, the introduction of local attention decreases the inference time by $2.96\%$ compared to the only global method. Visual comparions are demonstrated in Fig. 8, where local attention allows to capture region-specific blurred patterns and restore images with more visual details.
@@
Network Component PSNR Params (M) FLOPs (G)
Feature fusion AFF [6]$+$SFF $31.88$ 9.123 100.31
LDFF(SA$=$True)$+$SFF $31.88$ 9.096 99.71
LDFF(SA$=$True)$+$SGFF $31.89$ 9.079 99.28
LDFF$+$SGFF $31.89$ 8.995 97.13
LDFF$+$LDFF 31.96 9.082 99.41
Table 5: Influence of LCCL and DMFN. In the LCCL group, the positions of LCCL and impacts of different global pooling methods are investigated.
@@
LCCL CMSA w/o LCCL $31.98/0.953$ 99.41 19.7
LCCL before SA$+$GAP $31.99/0.954$ 99.46 22.6
LCCL after SA$+$GAP $32.01/0.954$ 99.46 22.2
@@
Effect of LDFF. Several fusion units are illustrated in Fig. 3. In Tab. 4, replacing AFF [6] by LDFF (enhanced with SA) yields the same quality scores with $27$ thousands fewer parameters and $0.6$ G fewer FLOPs. Tab. 4 shows that the self-attention mechanism does not improve the metric scores but brings about additional computational cost in LDFF. It also presents that applying LDFF to both multi-scale and same-scale feature fusion provides an overall gain of $0.08$ dB and improved efficiency over the baseline. Moreover, removing the gating mechanism from LDFF results in a noticeable $0.01$dB performance drop in an architecture without any activation in feed-forward networks. The reason is that our non-linear LDFF with has a stronger ability to model and propagate complex data across DeblurDiNAT, of which the decoder path contains no activation.
Effect of LCCL. We investigate how different positions of cross-channel learner (LCCL) in CMSA and various pooling strategies used in LCCL affect deblurring performance. Tab. 5 shows that, our parallel design obtains a $0.08$ dB gain and $2.26\%$ speedup over the strategy placing LCCL before the self-attention blocks. The global average pooling (GAP) in LCCL leads to $0.18$ dB boost over the global max pooling (GMP). Compared to the self-attention without channel modulation, our proposed CMSA improves PSNR by $0.09$ dB.
Effect of DMFN. Experimental data in Tab. 5 demonstrates that our straightforward DMFN brings a performance improvement of $0.07$ dB PSNR compared to GDFN [45] with gating mechanisms. DMFN also offers a $1.5\%$ speedup. Both of them share similar numbers of FLOPs.
Analysis on Transformer Blocks. We compare CMSA with self-attention mechanisms used in Restormer [45], Stripformer [38] and FFTformer [24], three SOTA Transformers designed for image deblurring. For a fair comparison, We replace both the self-attention CMSA and feed-forward network DMFN in DeblurDiNAT with their [45, 38, 24] corresponding ones. As reported in Tab. 6, ours outperforms the other attention methods in terms of PSNR and GPU memory consumption. Compared to SA employed in FFTformer [24], ours achieves $0.04$ dB higher PSNR, 2.45×2.45$\times$2.45 × faster speed and $49\%$ less GPU memory consumption.
Table 6: Comparison experiments among different Transformer blocks.
@@
CMSA 32.08 9.082 99.46 22.1 4.73